# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Appendix D Proof of Theorem 3
Proof of Lemma 7. Define a $\rho$-cover of $S^{d-1}$ as a set of $\mathcal{V}:=\{\mathbf{v}_{1},\dotsc,\mathbf{v}_{N}\}$ such that for any $\bm{\theta}\in S^{d-1}$, there exists some $\mathbf{v}_{i}$ such that $\langle\bm{\theta},\mathbf{v}_{i}\rangle\geq\rho$. The $\rho$-covering number of $S^{d-1}$ is
$N(\rho,S^{d-1}):=\inf\{N\in\mathbb{N}:\text{there exists a $\rho$-cover $\mathcal{V}$ of $S^{d-1}$ with size $N$}\}\,.$
By a simple argument we have $M(\rho,S^{d-1})\geq N(\rho,S^{d-1})$. Concretely, we construct a $\rho$-packing greedily by adding an element at each step which has inner product at most $\rho$ with all the previously selected elements, until it is no longer possible. This means that any point on $S^{d-1}$ has inner product larger than $\rho$ by some of the elements in the constructed set (otherwise it contradicts its maximality). Hence, we have a set that is both a $\rho$-cover and a $\rho$-packing of $S^{d-1}$, and by definition it results in $M(\rho,S^{d-1})\geq N(\rho,S^{d-1})$.
We next lower bound $N(\rho,S^{d-1})$ via a volumetric argument. Let $\mathcal{V}:=\{\mathbf{v}_{1},\dotsc,\mathbf{v}_{N}\}$ be a $\rho$-cover of $S^{d-1}$. For each element $\mathbf{v}_{i}\in\mathcal{V}$ we consider the cone around it with apex angle $\textup{cos}^{-1}(\rho)$. Its intersection with $S^{d-1}$ defines a spherical cap which we denote by $\mathcal{C}(\mathbf{v}_{i},\rho)$. Since $\mathcal{V}$ forms a $\rho$-cover of $S^{d-1}$, we have
$\textup{Vol}(S^{d-1})\leq\textup{Vol}(\cup_{i=1}^{N}\mathcal{C}(\mathbf{v}_{i},\rho))\leq\sum_{i=1}^{N}\textup{Vol}(\mathcal{C}(\mathbf{v}_{i},\rho))\,.$
We next use Lemma 2.2 from Ball et al., (1997) by which we have $\frac{\mathcal{C}(\mathbf{v}_{i},\rho)}{\textup{Vol}(S^{d-1})}\leq e^{-d\rho^{2}/2}$. Using this above, we get
@@
for any $\rho$-cover $\mathcal{V}$. Thus, we have $N(\rho,S^{d-1})\geq\exp(\frac{d\rho^{2}}{2})$, which completes the proof of the lemma.
Appendix E Proof of Theorem 4
Proof.
Recall the problem setting at the beginning of Section 4. Without loss of generality, we can prove the result for $\bm{\theta}^{\ast}={\mathbf{e}}_{1}$ and =$\mathcal{B}=\{{\mathbf{e}}_{1},{\mathbf{e}}_{2},\ldots,{\mathbf{e}}_{d}\}$ (just like the proof of Theorem 1). In that case, ${\mathrm{a}}^{(j)}={\mathrm{x}}^{(j)}=\langle{\mathbf{x}},{\mathbf{e}}_{j}\rangle$ for $j\in[d]$.
Let ${\mathrm{y}}=\text{sign}({\mathrm{x}}^{(1)})$ be the ground truth label of ${\mathbf{x}}$. Note that proving Theorem 4 is equivalent to proving the following statement with $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}\gamma^{2}}{12d}(1-2p)^{2}\Big{)}$. For a sample ${\mathbf{x}}\notin\mathcal{X}_{L}\cup\mathcal{X}_{U}$ such that $|\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}|\geq\frac{{16p_{2}}}{\mu}$, the following holds with a probability of at least $1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p_{2})\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$:
$\begin{cases}\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle\geq 0\text{ when }{\mathrm{y}}=1\text{ i.e., }{\mathrm{x}}^{(1)}\geq 0\\ \langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle<0\text{ when }{\mathrm{y}}=-1\text{ i.e., }{\mathrm{x}}^{(1)}<0.\end{cases}$ (100)
Also note that $\big{|}\textup{cos}\big{(}\beta({\mathbf{x}})\big{)}\big{|}=\frac{|{\mathrm{x}}^{(1)}|}{\|{\mathbf{x}}\|}$ in our choice of basis for the proof.
Just like the proof of Theorem 1, we will prove the above for the case of ${\mathrm{y}}=1$ or equivalently $\mathrm{x}^{(1)}\geq 0$, and the result for the other case (i.e., ${\mathrm{y}}=-1$ or equivalently $\mathrm{x}^{(1)}<0$) can be derived similarly.
Recall that:
$\widehat{\bm{\theta}}_{1}:=\frac{1}{\sigma^{2}}\Bigg{(}\frac{1}{n}\sum_{j=1}^{n}\widetilde{\mathrm{y}}_{j}^{\prime}{\mathbf{x}}_{j}^{\prime}\Bigg{)}.$ (101)
Note that:
$\widetilde{\mathrm{y}}_{j}^{\prime}=\begin{cases}\mathrm{y}_{j}^{\prime}\text{w.p. }\geq\alpha({\mathbf{x}}_{j}^{\prime})\\ -\mathrm{y}_{j}^{\prime}\text{ w.p. }\leq 1-\alpha({\mathbf{x}}_{j}^{\prime}),\end{cases}$ (102)
where $\alpha({\mathbf{x}})=1-\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}})\big{)}\Big{)}$ is as defined in Theorem 1.
We have:
@@ Appendix E Proof of Theorem 4
$\langle{\mathbf{x}},\widehat{\bm{\theta}}_{1}\rangle=\frac{1}{\sigma^{2}}\Bigg{\{}\frac{1}{n}\sum_{k=1}^{n}\widetilde{\mathrm{y}}_{k}^{\prime}\langle{\mathbf{x}}_{k}^{\prime},{\mathbf{x}}\rangle\Bigg{\}}.$ (103)
Since $\frac{1}{n\sigma^{2}}$ is a constant independent of $\mathcal{T}$ and $\mathcal{T}_{2}$, we will analyze
$\varepsilon^{\prime}=\sum_{k=1}^{n}\widetilde{\mathrm{y}}_{k}^{\prime}\langle{\mathbf{x}}_{k}^{\prime},{\mathbf{x}}\rangle.$ (104)
Using the Chernoff bound, we have for any $r<0$:
$\mathbb{P}\big{(}\varepsilon^{\prime}\leq 0\big{)}$ $=\mathbb{P}\big{(}\exp\big{(}r\varepsilon^{\prime}\big{)}\geq 1\big{)}\leq\mathbb{E}\Big{[}\exp\big{(}r\varepsilon^{\prime}\big{)}\Big{]}.$ (105)
For $k\in[n]$, let $\varepsilon_{k}^{\prime}:=\widetilde{\mathrm{y}}_{k}^{\prime}\langle{\mathbf{x}}_{k}^{\prime},{\mathbf{x}}\rangle$; thus, $\varepsilon^{\prime}=\sum_{k\in[n]}\varepsilon_{k}^{\prime}$. For any $r\in\mathbb{R}$, we have:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon^{\prime}\big{)}\Big{]}=\mathbb{E}\Big{[}\exp\big{(}r\sum_{k\in[n]}\varepsilon_{k}^{\prime}\big{)}\Big{]}=\mathbb{E}\Big{[}\prod_{k\in[n]}\exp\big{(}r\varepsilon_{k}^{\prime}\big{)}\Big{]}=\Big{(}\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}^{\prime}\big{)}\Big{]}\Big{)}^{n},$ (106)
where the last step follows because of the i.i.d. nature of the data. Now:
$\mathbb{E}\Big{[}\exp\big{(}r\varepsilon_{1}^{\prime}\big{)}\Big{]}$ $=\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}.$ (107)
Taking expectation w.r.t. the randomness in $\widetilde{\mathrm{y}}_{1}^{\prime}$ first, we get:
$\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}=\mathbb{E}\Bigg{[}\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\big{|}\widetilde{\mathrm{y}}_{1}^{\prime}={\mathrm{y}}_{1}^{\prime}\Big{]}\underbrace{\mathbb{P}\Big{(}\widetilde{\mathrm{y}}_{1}^{\prime}={\mathrm{y}}_{1}^{\prime}\Big{)}}_{=1-\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}}\\ +\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\big{|}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\Big{]}\mathbb{P}\Big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\Big{)}\Bigg{]}.$ (108)
$\mathbb{E}\Big{[}\exp\big{(}r\widetilde{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}=\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}+\underbrace{\mathbb{E}\Bigg{[}\Big{(}\exp\big{(}-r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{)}\mathbb{P}\Big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\Big{)}\Bigg{]}}_{:=\text{(I)}}.$ (109)
We already evaluated $\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}$ in the proof of Theorem 1; see eq. 20. It was:
@@
$\mathbb{E}\Big{[}\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{]}\leq\exp\big{(}\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\Bigg{(}\frac{(\sigma r\|{\mathbf{x}}\|)^{2}}{2}\Bigg{)}.$ (110)
[…]
From eq. 102, we have $\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}\leq 1-\alpha({\mathbf{x}}_{1}^{\prime})=\exp\Big{(}-\frac{n\mu^{2}}{4}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}}_{1}^{\prime})\big{)}\Big{)}$. Thus:
$\mathbb{E}\Bigg{[}\Big{(}\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}\Big{)}^{2}\Bigg{]}\leq\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{2}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}}_{1}^{\prime})\big{)}\Bigg{)}\Bigg{]}.$ (112)
Similar to eq. 64 in the proof of Theorem 2, it can be shown that:
$\mathbb{E}\Bigg{[}\exp\Bigg{(}-\frac{n\mu^{2}}{2}(1-2p)^{2}\textup{cos}^{2}\big{(}\beta({\mathbf{x}}_{1}^{\prime})\big{)}\Bigg{)}\Bigg{]}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{6d}(1-2p)^{2}\gamma^{2}\Bigg{)},$ (113)
and so:
$\mathbb{E}\Bigg{[}\Big{(}\mathbb{P}\big{(}\widetilde{\mathrm{y}}_{1}^{\prime}=-{\mathrm{y}}_{1}^{\prime}\big{)}\Big{)}^{2}\Bigg{]}\leq\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{6d}(1-2p)^{2}\gamma^{2}\Bigg{)}.$ (114)
Next,
$\mathbb{E}\Bigg{[}\Big{(}\exp\big{(}-r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{)}^{2}\Bigg{]}$ $=\mathbb{E}\Big{[}\exp\big{(}-2r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}+\exp\big{(}2r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-2\Big{]}.$ (115)
Using eq. 20 and eq. 21 in the proof of Theorem 1 above, we get:
$\mathbb{E}\Bigg{[}\Big{(}\exp\big{(}-r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}-\exp\big{(}r{\mathrm{y}}_{1}^{\prime}\langle{\mathbf{x}}_{1}^{\prime},{\mathbf{x}}\rangle\big{)}\Big{)}^{2}\Bigg{]}\leq\\ \Big{(}\underbrace{\exp\big{(}-2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}+\exp\big{(}2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}}_{=2\text{cosh}(2\mu\sigma r{\mathrm{x}}^{(1)})}\Big{)}\exp\Big{(}2{(\sigma r\|{\mathbf{x}}\|)^{2}}\Big{)}-2.$ (116)
For brevity, let $h(r):=2\text{cosh}\big{(}2\mu\sigma r{\mathrm{x}}^{(1)}\big{)}\exp\big{(}2{(\sigma r\|{\mathbf{x}}\|)^{2}}\big{)}-2$. Plugging in eq. 114 and eq. 116 into eq. 111 yields:
(I) $\leq\sqrt{h(r)}\sqrt{\exp\Bigg{(}-\frac{9(d-1)}{20}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{6d}(1-2p)^{2}\gamma^{2}\Bigg{)}}$
$\leq\sqrt{h(r)}\underbrace{\Bigg{(}\exp\Bigg{(}-\frac{9(d-1)}{40}\Bigg{)}+\exp\Bigg{(}-\frac{n\mu^{2}}{12d}(1-2p)^{2}\gamma^{2}\Bigg{)}\Bigg{)}}_{=p_{2}},$ (117)
where the last step is obtained using the fact that $\sqrt{a+b}\leq\sqrt{a}+\sqrt{b}$ for any $a,b\geq 0$. Also, let us put $p_{2}=\exp\Big{(}-\frac{9(d-1)}{40}\Big{)}+\exp\Big{(}-\frac{n\mu^{2}}{12d}(1-2p)^{2}\gamma^{2}\Big{)}$. Here we consider the case of $n$ and $d$ being large enough so that $p_{2}\leq\frac{1}{2}$. Now plugging in eq. 110 and eq. 117 into eq. 109 gives us:

### v2

@@ B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
Proof of Lemma D.3. Define a $\rho$-cover of $\mathbb{S}^{d-1}$ as a set of $\mathcal{V}:=\{\bm{v}_{1},\dotsc,\bm{v}_{N}\}$ such that for any $\bm{\theta}\in\mathbb{S}^{d-1}$, there exists some $\bm{v}_{i}$ such that $\langle\bm{\theta},\bm{v}_{i}\rangle\geq\rho$. The $\rho$-covering number of $\mathbb{S}^{d-1}$ is
$N(\rho,\mathbb{S}^{d-1}):=\inf\{N\in\mathbb{N}:\text{there exists a $\rho$-cover $\mathcal{V}$ of $\mathbb{S}^{d-1}$ with size $N$}\}\,.$
By a simple argument we have $M(\rho,\mathbb{S}^{d-1})\geq N(\rho,\mathbb{S}^{d-1})$. Concretely, we construct a $\rho$-packing greedily by adding an element at each step which has inner product at most $\rho$ with all the previously selected elements, until it is no longer possible. This means that any point on $\mathbb{S}^{d-1}$ has inner product larger than $\rho$ by some of the elements in the constructed set (otherwise it contradicts its maximality). Hence, we have a set that is both a $\rho$-cover and a $\rho$-packing of $\mathbb{S}^{d-1}$, and by definition it results in $M(\rho,\mathbb{S}^{d-1})\geq N(\rho,\mathbb{S}^{d-1})$.
We next lower bound $N(\rho,\mathbb{S}^{d-1})$ via a volumetric argument. Let $\mathcal{V}:=\{\bm{v}_{1},\dotsc,\bm{v}_{N}\}$ be a $\rho$-cover of $\mathbb{S}^{d-1}$. For each element $\bm{v}_{i}\in\mathcal{V}$ we consider the cone around it with apex angle $\textup{cos}^{-1}(\rho)$. Its intersection with $\mathbb{S}^{d-1}$ defines a spherical cap which we denote by $\mathcal{C}(\bm{v}_{i},\rho)$. Since $\mathcal{V}$ forms a $\rho$-cover of $\mathbb{S}^{d-1}$, we have
$\textup{Vol}(\mathbb{S}^{d-1})\leq\textup{Vol}(\cup_{i=1}^{N}\mathcal{C}(\bm{v}_{i},\rho))\leq\sum_{i=1}^{N}\textup{Vol}(\mathcal{C}(\bm{v}_{i},\rho))\,.$
We next use Lemma 2.2 from Ball et al., (1997) by which we have $\frac{\mathcal{C}(\bm{v}_{i},\rho)}{\textup{Vol}(\mathbb{S}^{d-1})}\leq e^{-d\rho^{2}/2}$. Using this above, we get
@@
for any $\rho$-cover $\mathcal{V}$. Thus, we have $N(\rho,\mathbb{S}^{d-1})\geq\exp(\frac{d\rho^{2}}{2})$, which completes the proof of the lemma.
Appendix E Proof of Theorem 4.8
Similar to the proof of Theorem 4.1, we want to upper bound $\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)$. Plugging in ${\widehat{\bm{\theta}}}_{1}$ from (4.7) we have
$y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle=\frac{y}{n}\langle\bm{x},\sum_{i}\widetilde{y}_{i}\bm{x}_{i}\rangle=\frac{y}{n}\langle\bm{x},\sum_{i}{\rm sign}(\langle\bm{x}_{i},{\widehat{\bm{\theta}}}_{0}\rangle)\bm{x}_{i}\rangle$
A major complication is that ${\widehat{\bm{\theta}}}_{0}$ depends on all of the data points in the training set. Expanding ${\widehat{\bm{\theta}}}_{0}$ we have
${\widehat{\bm{\theta}}}_{0}=\frac{1}{n}\sum_{i}\widehat{y}_{i}\bm{x}_{i}=\frac{1}{n}\sum_{i}\xi_{i}y_{i}\bm{x}_{i}=\frac{1}{n}\left((\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\sum_{i}\xi_{i}y_{i}{\bm{z}}_{i}\right)$ (E.1)
Here is the outline of our proof:
  • For every $\ell\in[n]$ we define the label
    $\widetilde{y}_{\ell}^{\prime}={\rm sign}\left(\langle\bm{x}_{\ell},(\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\,.$ (E.2)
  • We define the event :=$\mathcal{E}:=\{\forall\ell:\widetilde{y}^{\prime}_{\ell}={\rm sign}(\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle)\}$.
  • We write
    $\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0)$ $=\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E})+\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E}^{c})$
    $\leq\mathbb{P}(y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle<0;\mathcal{E})+\mathbb{P}(\mathcal{E}^{c})\,.$ (E.3)
  • We bound each of the term on the right-hand side separately. Note that under the event $\mathcal{E}$, we have
    $y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle=\frac{y}{n}\langle\bm{x},\sum_{\ell}\widetilde{y}_{\ell}^{\prime}\bm{x}_{\ell}\rangle.$
    We will control the probability of the RHS above being negative using concentration results on the sum of i.i.d. subgaussian random variables.
Next we get into details of this proof sketch. We start by bounding $\mathbb{P}(\mathcal{E}^{c})$.
Lemma E.1
Fix $\ell\in[n]$. Suppose that $nd\geq\frac{\gamma^{4}}{\lambda_{\max}^{2}}$ and define $p^{\prime}:=(1+\frac{3\gamma^{4}}{8\lambda_{\max}^{2}nd})p$. We then have
@@
$\mathbb{P}\left(\widetilde{y}_{\ell}^{\prime}\langle\bm{x}_{\ell},{\widehat{\bm{\theta}}}_{0}\rangle<0\right)\leq\frac{1}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$
Using Lemma E.1 along with union bounding over $\ell\in[n]$ (note that the events are dependent), we get
$\mathbb{P}(\mathcal{E}^{c})\leq\frac{n}{2}\left(\exp\left(-\frac{\gamma^{4}}{8\lambda_{\max}^{2}}(1-2p^{\prime})\frac{n}{d}\right)+e^{-d/16}\right)e^{d/n}\,.$ (E.4)
Next we note that under the event $\mathcal{E}$, we have:
$y\langle\bm{x},{\widehat{\bm{\theta}}}_{1}\rangle=\frac{y}{n}\langle\bm{x},\sum_{\ell}\widetilde{y}_{\ell}^{\prime}\bm{x}_{\ell}\rangle=\frac{1}{n}\sum_{\ell\in[n]}y{\rm sign}\left(\langle\bm{x}_{\ell},(\sum_{i}\xi_{i}(1+u_{i})){\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\langle\bm{x},\bm{x}_{\ell}\rangle.$ (E.5)
We define the shorthand $\beta:=\sum_{i}\xi_{i}(1+u_{i})$ and condition on $\beta$. Then (E.5) will be sum of i.i.d. subgaussian variables
$T_{\ell}:=y{\rm sign}\left(\langle\bm{x}_{\ell},\beta{\bm{{\mu}}}+\bm{\Sigma}^{1/2}\xi_{\ell}y_{\ell}{\bm{z}}_{\ell}\rangle\right)\langle\bm{x},\bm{x}_{\ell}\rangle\,,\quad\ell\in[n]\,.$ (E.6)
We continue by first characterizing its expectation.
Lemma E.2
Assuming $\frac{\beta\gamma^{2}}{2\lambda_{\max}d}<1$, for $\ell\in[n]$ we have
$\operatorname{\mathbb{E}}\left[T_{\ell}|\beta\right]\geq(1-2q)\langle y\bm{x},{\bm{{\mu}}}\rangle>0\,,$
where $q=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)$.
Recall that $\bm{x}_{\ell}=y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}$ and therefore
$\|T_{\ell}-\operatorname{\mathbb{E}}[T_{\ell}|\beta]\|_{\psi_{2}}\leq 2\|T_{\ell}\|_{\psi_{2}}$ $=2\|\langle\bm{x},\bm{x}_{\ell}\rangle\|_{\psi_{2}}$
$=2\|\langle\bm{x},y_{\ell}(1+u_{\ell}){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}_{\ell}\rangle\|_{\psi_{2}}$
$\leq 4|\langle\bm{x},{\bm{{\mu}}}\rangle|+2\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}\,.$
where we used the assumption $\|u_{\ell}\|_{\psi_{2}}=1$ and the fact that $\|{\bm{z}}\|_{\psi_{2}}=1$.
Next by using Hoeffding-type inequality for sum of sub-gaussian random variables (see e.g (Vershynin,, 2010, Proposition 5.1)) we have for every $t\geq 0$,
$\mathbb{P}\left(\Big{|}\sum_{\ell}T_{\ell}-\sum_{\ell}\operatorname{\mathbb{E}}[T_{\ell}|\beta]\Big{|}\geq t\Big{|}\beta\right)\leq 2\exp\left(-\frac{t^{2}}{64n(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\,.$
Therefore,
$\mathbb{P}(\sum_{\ell}T_{\ell}<0\;|\beta)$ $\leq\mathbb{P}\left(|\sum_{\ell}T_{\ell}-\sum_{\ell}\operatorname{\mathbb{E}}[T_{\ell}|\beta]|\geq\sum_{\ell}\operatorname{\mathbb{E}}[T_{\ell}|\beta]\;\Big{|}\;\beta\right)$
$\leq 2\exp\left(-\frac{n(\operatorname{\mathbb{E}}[T_{\ell}|\beta])^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)$
$\leq 2\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\,,$ (E.7)
where the last step follows from Lemma E.2.
Our next step is to take expectation of the above with respect to $\beta$. Before proceeding we define the following event:
$\mathcal{E}_{0}:=\left\{\beta\geq n(1-2p)/2\right\}\,.$ (E.8)
Note that by Hoeffding-type inequality for sub-gaussian random variables and given that $\operatorname{\mathbb{E}}[\xi_{i}]=1-2p$, we have
$\mathbb{P}\left(\Big{|}\sum_{i}\xi_{i}(1+u_{i})-n(1-2p)(1+\operatorname{\mathbb{E}}[u_{i}])\Big{|}\geq t\right)\leq 2\exp\left(-\frac{t^{2}}{8n}\right)\,.$
By choosing $t=n(1-2p)/2$, and using the fact that $u_{i}>0$ for all $i$ we get
$\mathbb{P}\left(\sum_{i}\xi_{i}(1+u_{i})<n(1-2p)/2\right)\leq 2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
Hence,
$\mathbb{P}(\mathcal{E}_{0}^{c})\leq 2\exp\left(-\frac{n}{32}(1-2p)^{2}\right).$
We now continue by taking expectation of (E.7) with respect to $\beta$.
@@
$\mathbb{P}(\sum_{\ell}T_{\ell}<0)$ $\leq 2\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\right]$
$=2\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)(\mathbb{1}_{\mathcal{E}_{0}}+\mathbb{1}_{\mathcal{E}_{0}^{c}})\right]$
$\leq 2\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2q)^{2}\langle y\bm{x},{\bm{{\mu}}}\rangle^{2}}{64(\langle\bm{x},{\bm{{\mu}}}\rangle^{2}+\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2})}\right)\mathbb{1}_{\mathcal{E}_{0}}\right]+2\mathbb{P}(\mathcal{E}_{0}^{c})\,.$ (E.9)
On the event $\mathcal{E}_{0}$ we have
$\frac{\beta\gamma^{2}}{2\lambda_{\max}d}\geq\frac{n(1-2p)\gamma^{2}}{4\lambda_{\max}d}>1,$
by our requirement in the theorem statement. Therefore,
$q=\exp\left(-\frac{\beta\gamma^{2}}{20\lambda_{\max}}\right)\leq\exp\left(-\frac{n(1-2p)\gamma^{2}}{40\lambda_{\max}}\right):=q^{\prime}\,.$
Since $01-2q\geq 1-2q^{\prime}>$ 0 this implies that